# Model Diffusion for Certifiable Few-shot Transfer Learning

### v1

@@ 6 Conclusion
We have introduced a novel Sample-Then-Evaluate approach to transfer learning. Our STEEL is designed to facilitate non-trivial performance guarantees, even in the low-shot regime, through use of a discrete hypothesis space. Our results instantiating the bound demonstrate that for both LLM and visual transfer learning STEEL performs comparably to alternatives it terms of test accuracy, while being dramatically better in terms of ability to provably guarantee this performance level.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, and we mention a non-exhaustive concerns here. First, for LLM experiments, please note that the diffusion model is trained on adapters trained on user-specific data. We can use the diffusion model to interpolate between adapters to potentially avoid distributing the private model zoo among clients, but this does have potential for data leakage by memorization (Staab et al., 2024). It is worth investigating privacy preserving methods for such concerns (Miranda et al., 2024).
On the vision side of our experiments, despite our benchmark being on fine-grained classification tasks of publicly available datasets, we necessitate the reminder of ethical concerns in computer vision as our method is easily applicable across domains and applications (Waelen, 2023).
@@ A.2 Bound with Parameter-Level Quantization
In (Lotfi et al., 2024), they proposed a non-vacuous bound for the LLM based on the model parameter quantization (e.g., fixed-length floating point machine representation). There are several differences to our approach:
@@ B.1 Diffusion Architecture and Training Recipe
The time embedding is generated from the diffusion timestep using sinsusoidal embeddings as per the original DDPM model (Ho et al., 2020). The dimensionality of the sinusoidal embedding is equal to the Diffusion MLP hidden dimension. Subsequently, the embedding is transformed using a two-layer MLP with first layer expanding the dimension to 4×4$\times$4 × the network’s hidden dimension and the second layer downscaling again to original hidden dimension. For example, on the vision experiments, the time embedding network has hidden dimension of $4098\times 4$. Finally, to condition the time embedding computed by the two-layer MLP time embedding network for each layer, we apply a different linear transformation per diffusion hidden layer.
We train the diffusion model for 30K epochs for all experiments with a batch size of 1,024. We use the LAMB optimizer (You et al., 2020) with a learning rate of 0.01. For vision experiments, we found that we can continue improving performance if we continue training for a second stage of 10K more epochs. For the second stage, we use a one-cycle learning rate scheduler (Smith & Topin, 2018) with default hyperparameters (as found in pytorch). The maximum learning rate starts from 0.0004 reaching a maximum of 0.001 over a 1000 steps. We keep an exponential-moving average of the network weights throughout training with a decaying rate of 0.9999.
@@ B.2 Flan-T5 + LaMP Hyperparameters
For training the base model across all datasets, we use LaMP’s original recipe (Salemi et al., 2024). We use a batch size of 64, AdamW optimizer with a learning rate of 5e-5 and weight decay of size 0.0001. We use a linear warmup for the learning rate over 5% of the total number of training steps.
@@
Finally, we transform the SGD number of epochs to MeZO. The authors used 32$\times$ as many epochs as SGD in the original paper (Malladi et al., 2023). This translates to 640, 320, 320 epochs for LaMP-2, LaMP-3, and LaMP-5. We tested the three learning rates proposed to search over by the authors. We fixed a learning rate of 1e-3 across all datasets because we consistently found 1e-4 to not learn and 1e-2 to be unstable.
For LaMP, we sample 10K LoRA-XS modules from the diffusion model. We use k-means clustering on the diffusion samples to produce N clusters where N is chosen as the minimum Silhouette score. We evaluate N between [2,150] inclusive. For each cluster, we find the medoid; the adapter closest to the centroid of the cluster. During evaluation, we choose a cluster and evaluate all adapters therein. From the cluster, we short-list the best 15 adapters using the Flan-T5 training loss. On the best 15 adapters, we use text generation to produce an answer with greedy sampling. Using the LaMP benchmark proposed model selection metric for each dataset, we select the “winning” adapter.
@@ B.3 CLIP + CoOP Hyperparameters
To build the model zoo, we used the authors original hyperparameters to train CoOP because we found them to work the best. This is SGD with a learning rate of 0.002. For Flowers-101, we train for 200 epochs. For DTD, FGVCAircraft and CUBirds, we trained for 300 epochs and found a One Cycle learning rate useful to stabilize training. These same hyperparameters were used to evaluate SGD on novel tasks. For BBPT, we use the authors default hyperparameters (Yu et al., 2023). We found that the method converges within 8000 “API call”. We attempted to run for a budget of 20K as our diffusion model offers but found that performance did not improve. Please note that the original authors reduce the dimensionality of the prompt using a small network because evolutionary optimization struggles in high-dimension. They reduce dimensionality to 512. Nevertheless, since we train only 2-tokens (dimensionality=1,024), then we do not use the small dimensionality reduction network.
@@ Appendix C Extra Vision Results
Dataset Zero Shot SGD BBPT Model Zoo Diffusion
@@
CUBirds 83.82% 81.77% 85.80% 86.00% 86.72%
DescribableTextures 67.29% 69.97% 74.10% 72.12% 73.97%
FGVCAircraft 47.44% 53.00% 55.05% 54.37% 55.40%
Flowers-101 81.14% 83.55% 80.40% 74.25% 76.75%
Avg 69.92% 72.07% 73.84% 71.69% 73.21%
@@
CUBirds 83.82 85.82% 86.92% 86.77% 86.50%
DescribableTextures 67.29 75.95% 76.57% 75.22% 76.17%
FGVCAircrafts 47.44 52.30% 57.90% 55.82% 57.30%
Flowers-101 81.14 86.92% 85.05% 77.17% 79.50%
Avg 69.92 75.25% 76.61% 73.75% 74.87%
@@
CUBirds 83.82 88.30% 88.42% 87.50% 87.47%
DescribableTextures 67.29 81.02% 77.90% 77.40% 79.82%
FGVCAircrafts 47.44 58.65% 60.55% 57.55% 60.10%
Flowers-101 81.14 91.95% 87.07% 79.82% 81.42%
Avg 69.92% 79.98% 78.49% 75.57% 77.21%
@@
CUBirds 83.82 89.75% 88.40% 87.42% 87.32%
DescribableTextures 67.29 85.07% 81.47% 79.90% 81.77%
FGVCAircrafts 47.44 62.07% 61.55% 58.77% 61.72%
Flowers-101 81.14 94.30% 88.67% 80.32% 82.52%
Avg 69.92% 82.80% 80.02% 76.61% 78.34%
@@
CUBirds 83.82 90.32% 89.27% 87.97% 88.40%
DescribableTextures 67.29 87.95% 83.20% 79.25% 81.50%
FGVCAircrafts 47.44 65.57% 62.37% 61.02% 61.37%
Flowers-101 81.14 95.90% 90.15% 82.92% 84.90%
Avg 69.92% 84.94% 81.25% 77.79% 79.04%

### v2

@@ 6 Conclusion
We have introduced a novel Sample-Then-Evaluate approach to transfer learning. Our STEEL is designed to facilitate non-trivial performance guarantees, even in the low-shot regime, through use of a discrete hypothesis space. Our results instantiating the bound demonstrate that for both LLM and visual transfer learning STEEL performs comparably to alternatives it terms of test accuracy, while being dramatically better in terms of ability to provably guarantee this performance level. Finally, based on our analysis from Section 5.3, we discuss future work in Appendix D.
@@ A.2 Bound with Parameter-Level Quantization
In [15], they proposed a non-vacuous bound for the LLM based on the model parameter quantization (e.g., fixed-length floating point machine representation). There are several differences to our approach:
@@
A.3 PAC-Bayes Meta Learning Bound
In [39] they proposed an extension of the PAC-Bayes bound for meta learning. Their meta learning algorithm aims to learn a distribution over the adaptation algorithms $A(S)\to Q(\theta)$ where $S$ is the support training data for a task and $Q(\theta)$ is the posterior distribution in the PAC-Bayes theorem. In their paper, any adaptation algorithm $A$ is allowable provided that $A$ internally makes use of a prior distributions in the PAC-Bayes theorem (e.g., the adaptation algorithms that try to minimize the PAC-Bayes bound), so that $A(S)$ becomes a function of the prior. And the meta learning model is a distribution $\rho$ that can sample the prior, and the meta learning amounts to learning the distribution $\rho$.
Although they provided a concrete example in Appendix B.2 in their paper, the described meta training involves an optimization of a loss function that contains an optimal value of another problem, which is typically attained by gradient descent. Just like MAML-type algorithms with a large number of inner gradient steps, this renders the application of their theorem to large-scale model/data impractical. So we propose a more practical reptile-like adaptation strategy, essentially admitting a closed-form inner optimal solution, that is practical in large model/data scenarios. Note that under our setting the meta learning PAC-Bayes bound is still applicable since they said the theorem works for any adaptation algorithms.
A.3.1 Concrete Cases with (Reptile-Like) Quadratic-Regularized Adaptation Algorithms
Let $\theta$ be the parameters of a neural network or PEFT parameters. Denote $\dim(\theta)$ by $d$. The adaptation algorithm $A(S)$ returns a PAC-Bayes posterior $Q(\theta)=\mathcal{N}(\mu,\Sigma)$ for some task support data $S$ as input. We consider diagonal $\Sigma$. The algorithm $A$ itself is parametrized by the PAC-Bayes prior $P_{1}(\theta)=\mathcal{N}(\mu_{1},\Sigma_{1})$ since we only deal with those $A$s that internally use $P_{1}$ as a part of the algorithm without dependency on any other things. One candidate for $A$ is the algorithm that minimizes the PAC-Bayes bound with prior $P_{1}$ with respect to $Q(\theta)$. However, as said earlier, this incurs MAML-like optimization that is impractical.
Our choice of $A$ is as follows: (Step-1) We first find $(\mu^{SGD},\Sigma^{SGD})=\arg\min_{\mathcal{N}(\theta;\mu,\Sigma)}\mathbb{E}[l(\theta;S)]$ by gradient descent. Technically, this involves evaluating $l(\theta;S)$ with reparametrized $\theta=\mu+\sqrt{\Sigma}z$ for $z\sim\mathcal{N}(0,I)$, where the gradients take: $\nabla_{\mu}l(\theta;S)=\nabla_{\theta}l(\theta;S)$ and $\nabla_{\Sigma}l(\theta;S)=\nabla_{\theta}l(\theta;S)\cdot 0.5\Sigma^{-0.5}z$ by the chain rule. So it is as tractable as SGD since we only need conventional backprop $\nabla_{\theta}l(\theta;S)$ to get gradients with respect to $(\mu,\Sigma)$. (Step-2) Solve the following quadratic optimization:
$(\mu^{*},\Sigma^{*})\ =\ \arg\min_{\mu,\Sigma}$ $||(\mu,\Sigma)-(\mu^{SGD},\Sigma^{SGD})||^{2}\ +\ \alpha\cdot||(\mu,\Sigma)-(\mu_{1},\Sigma_{1})||^{2}$ (8)
The intuition is that the adapted model $(\mu,\Sigma)$ has to be close to the fitted $(\mu^{SGD},\Sigma^{SGD})$ and also close to the prior $(\mu_{1},\Sigma_{1})$. The trade-off coefficient $\alpha$ is user’s choice, but when inspired by the PAC-Bayes bound objective, we can set $\alpha=1/\sqrt{|S|}$. The main benefit of this quadratic-regularized adaptation algorithm is that we have the closed-form solution:
$\mu^{*}(\mu_{1},S)$ $=\frac{1}{1+\alpha}\mu^{SGD}+\frac{\alpha}{1+\alpha}\mu_{1}$ (9)
$\Sigma^{*}(\Sigma_{1},S)$ $=\frac{1}{1+\alpha}\Sigma^{SGD}+\frac{\alpha}{1+\alpha}\Sigma_{1}$ (10)
Note that $(\mu^{*},\Sigma^{*})$ are functions of the prior $(\mu_{1},\Sigma_{1})$, and so we used the function notation in (910) together with dependency on $S$ although $\mu^{*}$ does not depend on $\Sigma_{1}$ and $\Sigma^{*}$ similarly. We treat $(\mu^{SGD},\Sigma^{SGD})$ as constant.
Next we define the meta prior ($\pi$) and posterior ($\rho$) distributions for the meta learning PAC-Bayes bound. As suggested in [39],
$\pi(A)=\mathcal{N}((\mu_{1},\Sigma_{1});0,\kappa_{\pi}^{2}I_{2d})$ (11)
$\rho(A)=\mathcal{N}((\mu_{1},\Sigma_{1});M_{1},\kappa_{\rho}^{2}I_{2d})$ (12)
where $M_{1}=(M^{\mu}_{1},M^{\Sigma}_{1})$ is the parameters for the meta posterior $\rho$. Here $\kappa_{\pi}$ and $\kappa_{\rho}$ are the fixed scales, and in their paper they used $\kappa_{\pi}=100$ and $\kappa_{\rho}=10^{-3}$. Finally, they assume delta meta prior and posterior in the complexity term: $\mathbb{P}(A)=\mathbb{Q}(A)=\delta_{P_{1}}$.
This leads to the following meta training objective:
$\min_{M_{1}}\ \hat{R}(\rho)+\underbrace{\sqrt{\frac{\textrm{KL}(\rho||\pi)+\log(4\sqrt{N}/\epsilon)}{2N}}}_{=:T_{1}}\ +\ \underbrace{\sqrt{\frac{\textrm{KL}(\rho||\pi)+N\hat{E}(\rho)+\log(8nN/\epsilon)+1}{2nN}}}_{=:T_{2}}$ (13)
where
$\hat{R}(\rho)$ $=\mathbb{E}_{A\sim\rho}\Bigg{[}\frac{1}{N}\sum_{i=1}^{N}l(A(S_{i});S_{i})\Bigg{]}$ (14)
$\hat{E}(\rho)$ $=\mathbb{E}_{A\sim\rho}\Bigg{[}\frac{1}{N}\sum_{i=1}^{N}\textrm{KL}(A(S_{i})||P_{1})\Bigg{]}$ (15)
$\textrm{KL}(\rho||\pi)$ $=\frac{2d\kappa_{\rho}^{2}+||M_{1}||^{2}}{2\kappa_{\pi}^{2}}-d+2d\log\frac{\kappa_{\pi}}{\kappa_{\rho}}$ (16)
Recall that $n$ is the support data size and $N$ is the number of meta training tasks.
Meta training procedure (stochastic approximation version). Since it is nearly infeasible to compute the objective for all $N$ tasks (for the purpose of one gradient update for $M_{1}$), we adopt the stochastic approximate optimization. That is, for each sampled task batch $B$ (out of $N$ tasks), we compute the followings (stochastic estimates of the objective and the gradient):
  • $\hat{T}_{0}\approx\hat{R}(\rho)$ where
    $\hat{T}_{0}=\frac{1}{|B|}\sum_{i\in B}l(\theta^{z_{i}};S_{i}),$ (17)
    $\theta^{z_{i}}=\mu^{*}(\mu_{1},S_{i})+\sqrt{\Sigma^{*}(\Sigma_{1},S_{i})}\cdot z_{i},\ \ z_{i}\sim\mathcal{N}(0,I_{d})$ (18)
    $\mu_{1}=M_{1}^{\mu}+\kappa_{\rho}\cdot z^{\mu},\ \ z^{\mu}\sim\mathcal{N}(0,I_{d})$ (19)
    $\Sigma_{1}=M_{1}^{\Sigma}+\kappa_{\rho}\cdot z^{\Sigma},\ \ z^{\Sigma}\mathcal{N}(0,I_{d})$ (20)
    Here $\mu^{*}(\mu_{1},S_{i})$ and $\Sigma^{*}(\mu_{1},S_{i})$ are determined by (910). Note that $\hat{T}_{0}$ is a function of $M_{1}$, and the gradient of $\hat{T}_{0}$ with respect to $M_{1}$ can be obtained from $\nabla_{\theta^{z}}\hat{T}_{0}$ by the chain rule similarly as described previously.
  • $T_{1}$ as it is since it is easy closed-form.
  • $\hat{T}_{2}\approx T_{2}$ where $\hat{T}_{2}$ only replaces $\hat{E}(\rho)$ by the following from $T_{2}$:
    $\hat{E}(\rho)$ $\to\frac{1}{|B|}\sum_{i\in B}\textrm{KL}(A(S_{i})\ ||\ P_{1})$ (21)
    $=\frac{1}{|B|}\sum_{i\in B}\textrm{KL}(\mathcal{N}(\mu^{*},\Sigma^{*})\ ||\ \mathcal{N}(\mu_{1},\Sigma_{1}))$ (22)
    This has a closed form from the Gaussian KL formula.
As the above steps allow us to compute the meta training objective and its gradient, we can update $M_{1}$, and move on to a next batch.
Test-time bound. Once we have meta-learned $M_{1}$, we can compute the error bound for a new task that comes with the support data $S^{*}$. In essence, we solve the following optimization problem:
$Q^{*}=\arg\min_{Q}\ \hat{R}(Q)+\sqrt{\frac{\textrm{KL}(Q||P_{1})+\log(8n/\epsilon)+1}{2n}}$ (23)
We parametrize $Q=\mathcal{N}(\mu,\Sigma)$ and solve it with respect to $(\mu,\Sigma)$. First, $\hat{R}(Q)$ can be approximated by $l(\theta^{z};S^{*})$ with $\theta^{z}=\mu+\sqrt{\Sigma}\cdot z$, $z\sim\mathcal{N}(0,I_{d})$. We can use the chain rule similarly for $\nabla_{\mu,\Sigma}\hat{R}(Q)$. In the KL term, $P_{1}$ has to be sampled from $\rho$ (i.e., our learned $M_{1}$). We can use $P_{1}=\mathcal{N}(M_{1}^{\mu}+\kappa_{\rho}\cdot z^{\mu},M_{1}^{\Sigma}+\kappa_{\rho}\cdot z^{\Sigma})$ and compute the KL using the Gaussian KL formula. Once optimized, the minimum value of (23) becomes the error bound for this test task which holds with probability at least $1-\epsilon$.
@@ B.1 Diffusion Architecture and Training Recipe
The time embedding is generated from the diffusion timestep using sinsusoidal embeddings as per the original DDPM model [8]. The dimensionality of the sinusoidal embedding is equal to the Diffusion MLP hidden dimension. Subsequently, the embedding is transformed using a two-layer MLP with first layer expanding the dimension to 4×4$\times$4 × the network’s hidden dimension and the second layer downscaling again to original hidden dimension. For example, on the vision experiments, the time embedding network has hidden dimension of $4098\times 4$. Finally, to condition the time embedding computed by the two-layer MLP time embedding network for each layer, we apply a different linear transformation per diffusion hidden layer.
We train the diffusion model for 30K epochs for all experiments with a batch size of 1,024. We use the LAMB optimizer [37] with a learning rate of 0.01. For vision experiments, we found that we can continue improving performance if we continue training for a second stage of 10K more epochs. For the second stage, we use a one-cycle learning rate scheduler [27] with default hyperparameters (as found in pytorch). The maximum learning rate starts from 0.0004 reaching a maximum of 0.001 over a 1000 steps. We keep an exponential-moving average of the network weights throughout training with a decaying rate of 0.9999.
@@ B.2 Flan-T5 + LaMP Hyperparameters
For training the base model across all datasets, we use LaMP’s original recipe [24]. We use a batch size of 64, AdamW optimizer with a learning rate of 5e-5 and weight decay of size 0.0001. We use a linear warmup for the learning rate over 5% of the total number of training steps.
@@
Finally, we transform the SGD number of epochs to MeZO. The authors used 32$\times$ as many epochs as SGD in the original paper [17]. This translates to 640, 320, 320 epochs for LaMP-2, LaMP-3, and LaMP-5. We tested the three learning rates proposed to search over by the authors. We fixed a learning rate of 1e-3 across all datasets because we consistently found 1e-4 to not learn and 1e-2 to be unstable.
For LaMP, we sample 10K LoRA-XS modules from the diffusion model. We use k-means clustering on the diffusion samples to produce N clusters where N is chosen as the minimum Silhouette score. We evaluate N between [2,150] inclusive. For each cluster, we find the medoid; the adapter closest to the centroid of the cluster. During evaluation, we choose a cluster and evaluate all adapters therein. From the cluster, we short-list the best 15 adapters using the Flan-T5 training loss. On the best 15 adapters, we use text generation to produce an answer with greedy sampling. Using the LaMP benchmark proposed model selection metric for each dataset, we select the “winning" adapter.
@@ B.2.2 Model Zoo Size
B.2.3 Compute Resources
All experiments were conducted on 10xA40 GPUs. The base model finetuning for generating task specific models took a couple of hours with a batch size of 64 using 2 GPUs. Time to train an adapter for one user using one GPUs varies from dataset to dataset and from user to user depending on their available dataset size. This takes as low as 10 seconds and up to 2 mins per user. Diffusion model training on the collected model zoo varied among datasets too since each epoch was of different total number of steps because of different dataset sizes. We used one A40 GPU per diffusion model and train using the recipe described earlier. Finally, downstream adaptation using STEEL varies between datasets. For LaMP-2, this was 50 A40 GPU hours and LaMP-5 was the slowest with 1200 A40 GPU hour since it requires text generation.
@@ B.3 CLIP + CoOP Hyperparameters
To build the model zoo, we used the authors original hyperparameters to train CoOP because we found them to work the best. This is SGD with a learning rate of 0.002. For Flowers-101, we train for 200 epochs. For DTD, FGVCAircraft and CUBirds, we trained for 300 epochs and found a One Cycle learning rate useful to stabilize training. These same hyperparameters were used to evaluate SGD on novel tasks. For BBPT, we use the authors default hyperparameters [38]. We found that the method converges within 8000 “API call". We attempted to run for a budget of 20K as our diffusion model offers but found that performance did not improve. Please note that the original authors reduce the dimensionality of the prompt using a small network because evolutionary optimization struggles in high-dimension. They reduce dimensionality to 512. Nevertheless, since we train only 2-tokens (dimensionality=1,024), then we do not use the small dimensionality reduction network.
Finally, for MetaPB, we train the upstream model for 20 epochs on our model zoo which takes 24 hours similarly to our diffusion model training. We use Adam with default learning rate of 0.01 for both upstream and downstream. Downstream adaptation was done for 300 epochs with same Adam optimizer setting. We train both means and variances. Variances are clipped to 0.1 for numerical stability.
@@
Compute Resources: For SGD, training one task upstream for model zoo collection or downstream for gradient descent adaptation takes 10-15 seconds. BBPT downstream takes 10-15 seconds as well. Training the diffusion model on a model zoo, since we fix model zoo size in CLIP experiments, was the same among all experiments. Each diffusion model took 24 A40 GPU hours to train. STEEL takes 30 second to 1 minute downstream for adapting to a new task. Finally, meta-training phase of MetaPB takes 24 hours similarly to our diffusion model training. MetaPB downstream takes 15 second per task too.
@@ Appendix C Extra Vision Results
C.1 Accuracy Results Across Shots
Dataset Zero Shot SGD BBPT Meta-PB Model Zoo STEEL
@@ C.1 Accuracy Results Across Shots
CUBirds 83.82% 81.77% 85.80% 88.25% 86.00% 86.72%
DescribableTextures 67.29% 69.97% 74.10% 78.55% 72.12% 73.97%
FGVCAircraft 47.44% 53.00% 55.05% 56.46% 54.37% 55.40%
Flowers-101 81.14% 83.55% 80.40% 54.20% 74.25% 76.75%
Avg 69.92% 72.07% 73.84% 69.37% 71.69% 73.21%
@@
CUBirds 83.82 85.82% 86.92% 88.40% 86.77% 86.50%
DescribableTextures 67.29 75.95% 76.57% 78.99% 75.22% 76.17%
FGVCAircrafts 47.44 52.30% 57.90% 58.31% 55.82% 57.30%
Flowers-101 81.14 86.92% 85.05% 61.25% 77.17% 79.50%
Avg 69.92 75.25% 76.61% 71.74% 73.75% 74.87%
@@
CUBirds 83.82 88.30% 88.42% 88.77% 87.50% 87.47%
DescribableTextures 67.29 81.02% 77.90% 78.85% 77.40% 79.82%
FGVCAircrafts 47.44 58.65% 60.55% 58.26% 57.55% 60.10%
Flowers-101 81.14 91.95% 87.07% 65.44% 79.82% 81.42%
Avg 69.92% 79.98% 78.49% 72.83% 75.57% 77.21%
@@
CUBirds 83.82 89.75% 88.40% 88.68% 87.42% 87.32%
DescribableTextures 67.29 85.07% 81.47% 79.96% 79.90% 81.77%
FGVCAircrafts 47.44 62.07% 61.55% 58.56% 58.77% 61.72%
Flowers-101 81.14 94.30% 88.67% 68.72% 80.32% 82.52%
Avg 69.92% 82.80% 80.02% 73.98% 76.61% 78.34%
@@
CUBirds 83.82 90.32% 89.27% 89.24% 87.97% 88.40%
DescribableTextures 67.29 87.95% 83.20% 81.12% 79.25% 81.50%
FGVCAircrafts 47.44 65.57% 62.37% 59.84% 61.02% 61.37%
Flowers-101 81.14 95.90% 90.15% 71.23% 82.92% 84.90%
Avg 69.92% 84.94% 81.25% 75.36% 77.79% 79.04%
@@
C.2 Certifying Zero-Shot Performance
Median Bound Flowers CUB DTD FGVVCA
STEEL 0.40 (+0.03) 0.36 (-0.06) 0.42 (+0.11) 0.61 (+0.10)
ZS CLIP 0.43 0.30 0.53 0.71
Table 4: Comparing median bounds between STEELand zero-shot CLIP given a 16-shots support set.
One might assume that zero-shot CLIP is a strong certified baseline because it is evaluated without any fine-tuning, and the support and query sets are drawn i.i.d. from the same distribution. This can lead to the impression that its support-set accuracy reflects test-time performance. However, these are still empirical estimates and do not account for uncertainty due to sampling. In contrast, our certificates explicitly bound the generalization gap from limited support data. To facilitate a fairer comparison, we compute a Langford-style test-set bound (Theorem 3.3 by Langford [14]) using zero-shot CLIP’s support accuracy from Table 3, treating it as a proxy for true performance. The resulting guarantees are included in the table below. While zero-shot CLIP performs well empirically, our STEEL method achieves tighter guarantees and often stronger accuracy in the few-shot setting.
Appendix D Limitations and Future Work
As discussed in the main paper (see Section 5.3), increasing support set size continues to tighten the certificate without causing underfitting. Beyond this, there remain several promising directions for further improving certificate strength. One avenue is to refine the structure of the hypothesis class. The current bound treats hypotheses as independent, but in practice, many may yield highly correlated predictions. Accounting for this redundancy, for instance, by discounting the contribution of similar models, could lead to tighter bounds. Another opportunity lies in the use of hybrid hypothesis classes. For example, treating each sample as a mean of a Gaussian mixture component and applying PAC-Bayesian fine-tuning may yield posterior distributions that better fit the data while maintaining low complexity. These strategies suggest that the certification framework presented here can be further enhanced by incorporating richer modeling and more expressive prior structures.
Appendix E Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, and we mention a non-exhaustive concerns here. First, for LLM experiments, please note that the diffusion model is trained on adapters trained on user-specific data. We can use the diffusion model to interpolate between adapters to potentially avoid distributing the private model zoo among clients, but this does have potential for data leakage by memorization [29]. It is worth investigating privacy preserving methods for such concerns [18].
On the vision side of our experiments, despite our benchmark being on fine-grained classification tasks of publicly available datasets, we necessitate the reminder of ethical concerns in computer vision as our method is easily applicable across domains and applications [33].